# Higher spins and Finsler geometry

### v1

@@
Finsler geometry is a natural generalization of (pseudo-)Riemannian geometry, where the line element is not the square root of a quadratic form but a more general homogeneous function. Parameterizing this in terms of symmetric tensors suggests a possible interpretation in terms of higher-spin fields. We will see here that, at linear level in these fields, the Finsler version of the Ricci tensor leads to the curved-space Fronsdal equation for all spins, plus a Stueckelberg-like coupling. Nonlinear terms can also be systematically analyzed, suggesting a possible interacting structure. No particular choice of spacetime dimension is needed. The Stueckelberg mechanism breaks gauge transformations to a redundancy that does not change the geometry. This is however not enough to eliminate non-transverse modes, at least for some versions of Finsler dynamics.
@@ 1 Introduction
The dynamics of fields with spin higher than two has an intricate structure. Their free equations of motion are already quite complicated when they have a mass [1], but simplifies in the massless case, where their equation of motion mimics that of spin two, using the so-called Fronsdal operator ${\mathcal{F}}$ [2]. Their interactions have a highly constrained dynamics; for reviews see for example [3, 4, 5, 6, 7]. It has long been surmised that they are important for theories of quantum gravity. String theory does have a tower of such fields, with masses related to the string tension. On the other hand, theories of Vasiliev type describe massless higher spins in (A)dS [8, 9, 10, 11, 12]. Their status as quantum theories is bolstered by their holographic interpretation [13, 14]. Higher spin fields and symmetries are expected more generally in holography, although most often in a broken version [15]. Some versions of these theories can indeed be even obtained as limits of string theory [16, 17].
@@
$F^{2}=2g_{2}+\frac{\phi_{4}}{g_{2}}+\frac{\phi_{6}}{g_{2}^{2}}\ldots\,,\qquad\phi_{s}\equiv\frac{1}{s!}\phi_{\mu_{1}\ldots\mu_{s}}\dot{x}^{\mu_{1}}\ldots\dot{x}^{\mu_{s}}\,.$ (1.1)
@@
A version of (1.1) with $s=2$ and one $s>2$ was considered in this context [23, (5.8)], and an approximate gauge transformation for the particle action was noticed. In [24] an analogue of Finsler geometry (without the homogeneity constraint) was used for W-gravity in low dimensions. Most relevant for us, (1.1) was found in [25] for $s=4$ and $g_{2}=\eta_{2}$ (flat space) to be related to the spin-4 Fronsdal equation [2].
@@
$\rho\sim R^{0}_{\alpha\beta}y^{\alpha}y^{\beta}+\frac{1}{2}{\mathcal{F}}(\delta F^{2})$ (1.3)
@@
This essential lack of gauge transformations (beyond the usual diffeomorphisms) is worrisome. We will see that the equations of motion for a Finsler analogue of GR, the simplest of which just reads $\rho=0$, have many perturbative solutions that are not transverse to the momentum, and thus would presumably create trouble upon quantization. These solutions also have many free parameters; this might give hope that a gauge transformation can indeed be defined for this system. Unfortunately I have found no such candidate, but I have not excluded it either.
There are of course many other possible problems with a theory of higher spins [3]. The Vasiliev theory manages to avoid them in a rather sophisticated fashion. In contrast, the idea of a Finsler higher spin theory should be regarded as moving here one of its first steps. Nevertheless, the present results make it worth exploring further in the future.
@@ 2.4 Flat space
We will now see that (2.21) inherits some of the nice properties of (2.19). In this subsection we work in flat space, $g_{\mu\nu}=\eta_{\mu\nu}$.
@@ 3.2 Connection
The second equality uses (3.7).
@@
This property is called almost compatibility with the metric of the Chern connection (3.13). Other connections exist with nice properties, the most notable being the Cartan connection, which also has components along the $y$ coordinates; it is metric-compatible but has torsion [30, Thm. 1.4.2].
@@ 4.1 Deformations
It can be shown [18, Thm. 3.3.1] that the geodesics of $\tilde{F}^{2}$ coincide with those of $F^{2}$ iff $\delta_{\nu}\tilde{F}=y^{\mu}\partial_{y^{\nu}}\delta_{\mu}\tilde{F}$; in this case $\tilde{G}^{\mu}=G^{\mu}+Py^{\mu}$, with $P=\tilde{F}^{-1}y^{\nu}\delta_{\nu}\tilde{F}$. The two Finsler $\tilde{F}$ and $F$ are said to be projectively equivalent.
@@ 4.2 Parametrization
An alternative to (4.8) would be obtained by applying the same logic directly to $F$, leading to $F=\sqrt{2g_{2}}+f_{3}/g_{2}+f_{4}(g_{2})^{3/2}+\ldots$. Adding the spin-one term to this parameterization leads to a generalization of the Randers choice mentioned in Sec. 3.1.
@@ 5.1 General expansion
We still raise and lower indices with the unperturbed (pseudo-)Riemannian metric, now denoted by $g^{0}_{\mu\nu}$ to avoid confusion with the fundamental tensor (3.4) of Finsler geometry. We expand the latter and its inverse as in (4.8), (4.12):
@@ 5.2 Spin-six nonlinear equation
This signals a cubic interaction among two $s=4$ and one $s=6$ fields. (Such an interaction is also present in Vasiliev theories [31].)
@@ 6.1 Equations of motion
(See [20, 11.3.1] and [35].) The measure is the product of the usual $\sqrt{g}\mathrm{d}^{D}x$ and of $\sqrt{g}\mathrm{d}^{D-1}\hat{y}$ along the fiber of the sphere tangent bundle $SM$. The $F^{2}$ is included to make the integrand of degree zero, so that it is defined on the fiber of that bundle. The variation of (6.2) gives
@@ 6.2 Broken Fronsdal gauge transformations
The appearance of the Fronsdal operator in (4.14) might give hope that gauge transformations of the form $\delta\phi_{s}=\mathrm{d}\lambda_{s-1}$ can be introduced in Finsler geometry, at least for a small perturbation around flat space $F^{2}=2\eta_{2}$ (where (2.28) is valid). It is not clear why the $F^{2}$ parameterization (4.8) should be invariant under such a transformation.121212[23, (5.9)] notes that the invariance of proper time $\int\mathrm{d}\tau F$ can be approximately restored by simultaneously changing the position $x^{\mu}(\tau)$ in a velocity-dependent fashion. In other words, the gauge transformations of Finsler geometry are only diffeomorphisms. A priori it might be possible, however, that $F^{2}$ changes in a way that does not modify the geometry in a physically relevant way. We would then declare such transformations to be additional gauge transformations. This might perhaps be similar to the notion of projective equivalence mentioned in Sec. 4.1.131313The infinitesimal counterpart of (4.1) is $\delta G^{\mu}=\delta Py^{\mu}$ with $\delta P=(2g_{2})^{-1}\mathrm{d}\delta F=(2g_{2}^{3})^{-1/2}\mathrm{d}\delta F^{2}$. From (4.3) we now find $\delta\rho=(1-D/2)\mathrm{d}\delta P$; so $\rho$ is not invariant under a projective equivalence.
@@
One might try to cancel the two terms in the parenthesis against each other. The structure looks similar to that of a Stückelberg mechanism, where a mass term locks together the gauge transformations of two fields of different spin. It is also reminiscent of the compensator fields in [26, 37, 38], where an equation of the type ${\mathcal{F}}(\phi_{s})=\mathrm{d}^{3}{\mathcal{H}}_{s}$ appears, with $\delta{\mathcal{H}}_{s}=\lambda^{\prime}_{s-1}$. [^14] In the AdS case an extra term is present because of (2.39); we will deal with this at the end. For theories of Vasiliev type, a certain boundary condition for the scalar can lead to a Higgs mechanism whose Goldstone field is a bound state of higher spin fields [39].
@@
We have concluded that we cannot set to zero the parenthesis in (6.5), or $\Lambda^{\prime}$ in (6.8). In other words, we have tried to find gauge transformations with an algebraic constraint (similar to the traceless constraint for Fronsdal), and we failed. We could still hope that the full (6.5) can somehow be made zero, by playing with the differential operators. However, $\mathrm{d}^{3}$ has a finite-dimensional kernel, consisting of quadratic polynomials in the $x^{\mu}$. Thus we also cannot find gauge transformations this way. [^15] We can apply a similar argument to projective equivalence to flat space. The variation of (4.5) gives $\delta G_{\mathrm{L}}^{\mu}=\frac{1}{2}\mathrm{d}^{2}\partial_{y_{\mu}}\Lambda$. From below (4.1) we find $\delta G^{\mu}=Py^{\mu}$. If we ignore the finite-dimensional kernel of $\mathrm{d}^{2}$, we are led to postulate $\partial_{y_{\mu}}\Lambda=y^{\mu}Q$, with $\mathrm{d}^{2}Q=2P$. This implies a “purely radial” dependence, $\Lambda=\Lambda(\eta_{2})$, and the homogeneity property lets us conclude $\Lambda=\lambda_{0}\sqrt{\eta_{2}}$, which would correspond to the gauge transformations of a spin-1 field, had we included it in (4.8). This is compatible with [18, Prop. 3.4.8]; notice however that many other Finsler metrics projectively equivalent to flat space are given in that book, not obtained by gauge transformations of Fronsdal type.
We now also analyze this possibility using the parameterization in terms of sum over spins. It is lengthier, but instructive, and also a good warm-up for the next subsection. We earlier concluded that we can always fix $\lambda^{\prime}_{s-1}=0$. Moreover we recall from Sec. 4.2 that $\phi_{s}^{\prime}=0$ can similarly be assumed without loss of generality. Since $(\mathrm{d}\lambda)^{\prime}=\mathrm{d}\lambda^{\prime}+2\partial_{\mu}\lambda^{\mu}$, we also conclude
@@
We can also try a similar analysis on a maximally symmetric space such AdS. Here is what changes. First, covariant derivatives don’t commute, and that complicates the trace $(\mathrm{d}^{3}\lambda)^{\prime}$, which is now $(6\mathrm{d}(2\nabla^{2}-\Delta)-2[\mathrm{d},\nabla^{2}])\lambda_{s-1}$. Second, now $\delta{\mathcal{F}}\neq 0$ even if $\lambda^{\prime}_{s-1}=0$, as we saw back in (2.39). Third, it makes sense to consider now the variant of (6.1) discussed below it, $\rho-kF^{2}=0$. This however can be reabsorbed in (2.39) by redefining $\Lambda$. All in all (6.12) is changed to
@@ 6.4 Discussion
While the Finsler Ricci $\rho$ seems important in view of its connection to the Fronsdal operator, another possibility is that it might have to be modified by additional terms. Even more radically, one might have to modify the central tenets of Finsler geometry altogether, as advocated in [24]. It might be helpful to try and recast in terms of Finsler geometry the existing higher spin theories such as those by Vasiliev [8, 9, 10, 11, 12], the simpler alternatives in three dimensions [42, 43], or the more recent self-dual theories [44, 45]. In any case, after finding a geometrical quantity that is invariant under a set of infinite transformations that extends diffeomorphism invariance, we would declare such transformations to be additional gauge transformations, identifying different Finsler geometries that are equivalent as far as physics is concerned. Time will tell whether any of this can be achieved.
@@ Acknowledgements
It’s a pleasure to thank S. Giombi, C. Hull, E. Minguzzi, C. Pfeifer, A. Sagnotti, A. Zaffaroni for very useful discussions and correspondence. My interest in Finsler geometry was rekindled by an ongoing collaboration with B. De Luca, N. De Ponti, A. Mondino, to whom I am very grateful. I would also like to thank the participants in the 2024 Cook’s Branch Workshop, to whom a preliminary version of these findings was presented. I am supported in part by INFN and by MIUR-PRIN contract 2022YZ5BA2.

### v2

@@
Finsler geometry is a natural generalization of (pseudo-)Riemannian geometry, where the line element is not the square root of a quadratic form but a more general homogeneous function. Parameterizing this in terms of symmetric tensors suggests a possible interpretation in terms of higher-spin fields. We will see here that, at linear level in these fields, the Finsler version of the Ricci tensor leads to the curved-space Fronsdal equation for all spins, plus a Stueckelberg-like coupling. Nonlinear terms can also be systematically analyzed, suggesting a possible interacting structure. No particular choice of spacetime dimension is needed. The Stueckelberg mechanism breaks gauge transformations to a redundancy that does not change the geometry. This creates a serious issue: non-transverse modes are not eliminated, at least for the versions of Finsler dynamics examined in this paper.
@@ 1 Introduction
The dynamics of fields with spin higher than two has an intricate structure. Their free equations of motion are already quite complicated when they have a mass [1], but simplifies in the massless case, where their equation of motion mimics that of spin two, using the so-called Fronsdal operator ${\mathcal{F}}$ [2]. Their interactions have a highly constrained dynamics; for reviews see for example [3, 4, 5, 6, 7]. It has long been surmised that they are important for theories of quantum gravity. String theory does have a tower of such fields, with masses related to the string tension. On the other hand, theories of Vasiliev type describe massless higher spins in (A)dS [8, 9, 10, 11, 12]. Their status as quantum theories is bolstered by their holographic interpretation [13, 14], although some degree of non-locality emerges [15, 16]. Higher spin fields and symmetries are expected more generally in holography, although most often in a broken version [17]. Some versions of these theories can indeed be even obtained as limits of string theory [18, 19]. Restrictions to local covariant massless higher spins in flat space were found in [20, 21, 22]; more recently, progress has been made using light-cone approach [23, 22].
@@
$F^{2}=2g_{2}+\frac{\phi_{4}}{g_{2}}+\frac{\phi_{6}}{g_{2}^{2}}+\ldots\,,\qquad\phi_{s}\equiv\frac{1}{s!}\phi_{\mu_{1}\ldots\mu_{s}}\dot{x}^{\mu_{1}}\ldots\dot{x}^{\mu_{s}}\,.$ (1.1)
@@
A version of (1.1) with $s=2$ and one $s>2$ was considered in this context [29, (5.8)], and an approximate gauge transformation for the particle action was noticed. This was further explored and generalized in [30, 31], in a line of research that led to a conformal theory of higher spins [32]. [^3] These references use a Hamiltonian point of view, which was recently further explored in [33, 34] in the context of self-dual higher spins [23]. I thank D. Ponomarev for insightful comments. In [35] an analogue of Finsler geometry (without the homogeneity constraint) was used for W-gravity in low dimensions. Most relevant for us, (1.1) was found in [36] for $s=4$ and $g_{2}=\eta_{2}$ (flat space) to be related to the spin-4 Fronsdal equation [2].
@@
$\rho\sim R^{0}_{\alpha\beta}y^{\alpha}y^{\beta}-\frac{1}{2}{\mathcal{F}}(\delta F^{2})$ (1.3)
@@
This essential lack of gauge transformations (beyond the usual diffeomorphisms) is worrisome. We will see that the equations of motion for a Finsler analogue of GR, the simplest of which just reads $\rho=0$, have many perturbative solutions that are not transverse to the momentum, and thus would presumably create trouble upon quantization. These solutions also have many free parameters; this might give hope that a gauge transformation can indeed be defined for this system. Unfortunately I have found no such candidate, but I have not excluded it either. These negative results are not entirely unexpected, in light of some of the no-go results mentioned above. The spirit of this paper is that reexamining an issue in a new light can lead to new insights. Overall, the idea of a Finsler higher spin theory faces significant challenges. Nevertheless, in my opinion the present results make it worth exploring further in the future.
@@ 2.3 Differential operators
We will see soon that this equation is invariant under a gauge transformation, but only in maximally symmetric spaces.
@@ 2.4 Flat space
We will now see that in flat space (2.21) inherits some of the nice properties of (2.19). So in this subsection $g_{\mu\nu}=\eta_{\mu\nu}$.
@@ 3.2 Connection
The second equality uses (3.7). Recall that we are using here the fundamental tensor (3.4) to lower indices.
@@
This property is called almost compatibility with the metric of the Chern connection (3.13). Other connections exist with nice properties, the most notable being the Cartan connection, which also has components along the $y$ coordinates; it is metric-compatible but has torsion [41, Thm. 1.4.2].
@@ 4.1 Deformations
It can be shown [24, Thm. 3.3.1] that the geodesics of $\tilde{F}^{2}$ coincide with those of $F^{2}$ iff $\delta_{\nu}\tilde{F}=y^{\mu}\partial_{y^{\nu}}\delta_{\mu}\tilde{F}$; in this case $\tilde{G}^{\mu}=G^{\mu}+Py^{\mu}$, with $P=\tilde{F}^{-1}y^{\nu}\delta_{\nu}\tilde{F}$. The two Finsler $\tilde{F}$ and $F$ are said to be projectively equivalent.
@@ 4.2 Parametrization
An alternative to (4.8) would be obtained by applying the same logic directly to $F$, leading to $F=\sqrt{2g_{2}}+f_{3}/g_{2}+f_{4}/g_{2}^{3/2}+\ldots$. Adding the spin-one term to this parameterization leads to a generalization of the Randers choice mentioned in Sec. 3.1.
@@ 5.1 General expansion
In this section we raise and lower indices with the unperturbed (pseudo-)Riemannian metric, now denoted by $g^{0}_{\mu\nu}$ to avoid confusion with the fundamental tensor (3.4) of Finsler geometry. We expand the latter and its inverse as in (4.8), (4.12):
@@ 5.2 Spin-six nonlinear equation
Formally this looks like a cubic interaction among two $s=4$ and one $s=6$ fields.
@@ 6 Challenges for Finsler dynamics
As commented in the introduction, there are strong arguments against a flat space local covariant theory of massless higher spins. So the challenges we will find are not unexpected. Nevertheless, we will start from flat space for simplicity. We will find later that the situation does not seem to improve much for AdS.
@@ 6.1 Equations of motion
(See [26, 11.3.1] and [45].) The measure is the product of the usual $\sqrt{g}\mathrm{d}^{D}x$ and of $\sqrt{g}\mathrm{d}^{D-1}\hat{y}$ along the fiber of the sphere tangent bundle $SM$. The $F^{2}$ is included to make the integrand of degree zero, so that it is defined on the fiber of that bundle. The variation of (6.2) gives
@@ 6.2 Broken Fronsdal gauge transformations
The appearance of the Fronsdal operator in (4.14) might give hope that gauge transformations of the form $\delta\phi_{s}=\mathrm{d}\lambda_{s-1}$ can be introduced in Finsler geometry, at least for a small perturbation around flat space $F^{2}=2\eta_{2}$ (where (2.28) is valid). It is not clear why the $F^{2}$ parameterization (4.8) should be invariant under such a transformation.131313[29, (5.9)] notes that the invariance of proper time $\int\mathrm{d}\tau F$ can be approximately restored by simultaneously changing the position $x^{\mu}(\tau)$ in a velocity-dependent fashion. In other words, the gauge transformations of Finsler geometry are only diffeomorphisms. A priori it might be possible, however, that $F^{2}$ changes in a way that does not modify the geometry in a physically relevant way. We would then declare such transformations to be additional gauge transformations. This might perhaps be similar to the notion of projective equivalence mentioned in Sec. 4.1.141414The infinitesimal counterpart of (4.1) is $\delta G^{\mu}=\delta Py^{\mu}$ with $\delta P=(2g_{2})^{-1}\mathrm{d}\delta F=(2g_{2}^{3})^{-1/2}\mathrm{d}\delta F^{2}$. From (4.3) we now find $\delta\rho=(1-D/2)\mathrm{d}\delta P$; so $\rho$ is not invariant under a projective equivalence.
@@
One might try to cancel the two terms in the parenthesis against each other. The structure looks similar to that of a Stückelberg mechanism, where a mass term locks together the gauge transformations of two fields of different spin. It is also reminiscent of the compensator fields in [37, 47, 48, 49], where an equation of the type ${\mathcal{F}}(\phi_{s})=\mathrm{d}^{3}{\mathcal{H}}_{s}$ appears, with $\delta{\mathcal{H}}_{s}=\lambda^{\prime}_{s-1}$. [^15] In the AdS case an extra term is present because of (2.39); we will deal with this at the end. For theories of Vasiliev type, a certain boundary condition for the scalar can lead to a Higgs mechanism whose Goldstone field is a bound state of higher spin fields [50].
@@
We have concluded that we cannot set to zero the parenthesis in (6.5), or $\Lambda^{\prime}$ in (6.8). In other words, we have tried to find gauge transformations with an algebraic constraint (similar to the traceless constraint for Fronsdal), and we failed. We could still hope that the full (6.5) can somehow be made zero, by playing with the differential operators. However, $\mathrm{d}^{3}$ has a finite-dimensional kernel, consisting of quadratic polynomials in the $x^{\mu}$. Thus we also cannot find gauge transformations this way. [^16] We can apply a similar argument to projective equivalence to flat space. The variation of (4.5) gives $\delta G_{\mathrm{L}}^{\mu}=\frac{1}{2}\mathrm{d}^{2}\partial_{y_{\mu}}\Lambda$. From below (4.1) we find $\delta G^{\mu}=Py^{\mu}$. If we ignore the finite-dimensional kernel of $\mathrm{d}^{2}$, we are led to postulate $\partial_{y_{\mu}}\Lambda=y^{\mu}Q$, with $\mathrm{d}^{2}Q=2P$. This implies a “purely radial” dependence, $\Lambda=\Lambda(\eta_{2})$, and the homogeneity property lets us conclude $\Lambda=\lambda_{0}\sqrt{\eta_{2}}$, which would correspond to the gauge transformations of a spin-1 field, had we included it in (4.8). This is compatible with [24, Prop. 3.4.8]; notice however that many other Finsler metrics projectively equivalent to flat space are given in that book, not obtained by gauge transformations of Fronsdal type.
We now also analyze this possibility using the parameterization in terms of sum over spins. It is lengthier, but instructive, and also a good warm-up for the next subsection. We earlier concluded that we can always fix $\lambda^{\prime}_{s-1}=0$. Moreover we recall from Sec. 4.2 that $\phi_{s}^{\prime}=0$ can similarly be assumed without loss of generality. Since $(\mathrm{d}\lambda)^{\prime}=\mathrm{d}\lambda^{\prime}+2\partial_{\mu}\lambda^{\mu}$, we also conclude (as in [51]) that
@@
We can also try a similar analysis on a maximally symmetric space such AdS. This a priori might look more promising, in light of the aforementioned no-go arguments. Here is what changes. First, covariant derivatives don’t commute, and that complicates the trace $(\mathrm{d}^{3}\lambda)^{\prime}$, which is now $(6\mathrm{d}(2\nabla^{2}-\Delta)-2[\mathrm{d},\nabla^{2}])\lambda_{s-1}$. Second, now $\delta{\mathcal{F}}\neq 0$ even if $\lambda^{\prime}_{s-1}=0$, as we saw back in (2.39). Third, it makes sense to consider now the variant of (6.1) discussed below it, $\rho-kF^{2}=0$. This however can be reabsorbed in (2.39) by redefining $\Lambda$. All in all (6.12) is changed to
@@ 6.4 Discussion
While the Finsler Ricci $\rho$ seems important in view of its connection to the Fronsdal operator, another possibility is that it might have to be modified by additional terms. Even more radically, one might have to modify the central tenets of Finsler geometry altogether, as advocated in [35]. In a similar spirit, one might try to adapt the Hamiltonian approach in [30, 31]. Its advantage is that higher-spin symmetries are manifest as canonical transformations. It is possible to define a so-called “coFinsler” distance in this way, and it would be interesting to see how the present results are modified in that language. One should however recall that the results in [30, 31] were partially negative (see also [33, 34]).
It might be helpful to try and recast in terms of Finsler geometry the existing higher spin theories such as those by Vasiliev [8, 9, 10, 11, 12], the simpler alternatives in three dimensions [54, 55], or the more recent self-dual theories [23, 22, 56, 57]. [^18] In a twistor-inspired approach to gravity, Finsler geometry appears rather naturally [58]. Finally, the higher spin symmetries might be realized around a solution where all the higher spin fields are non-zero, rather than around a (pseudo-)Riemannian geometry as attempted here. [^19] We thank M. Montero and C. Pfeifer for this suggestion.
In any case, after finding a geometrical quantity that is invariant under a set of infinite transformations that extends diffeomorphism invariance, we would declare such transformations to be additional gauge transformations, identifying different Finsler geometries that are equivalent as far as physics is concerned. Time will tell whether any of this can be achieved.
@@ Acknowledgements
It’s a pleasure to thank J. Calderón Infante, S. Giombi, C. Hull, L. Mason, E. Minguzzi, M. Montero, C. Pfeifer, D. Ponomarev, A. Sagnotti, A. Zaffaroni for very useful discussions and correspondence. My interest in Finsler geometry was rekindled by an ongoing collaboration with B. De Luca, N. De Ponti, A. Mondino, to whom I am very grateful. I would also like to thank the participants in the 2024 Cook’s Branch Workshop, to whom a preliminary version of these findings was presented. I am supported in part by INFN and by MIUR-PRIN contract 2022YZ5BA2.